### v2

# Cramer-Rao Bounds for Target Parameter Estimation in a Bi-Static IRS-Assisted Radar Configuration

Abstract
The use of Intelligent Reflective Surfaces (IRS) to assist communication and sensing has proven cost-effective in challenging scenarios. For sensing, IRS is shown to sense non-line-of-sight (NLOS) and stealth targets, albeit with significant loss due to the four-hop path model. Amongst the available IRS-assisted configurations, we consider a three-hop model in which the IRS redirects the scattered target response towards the mono-static radar. With the IRS spatially displaced from the radar, this configuration mimics a bi-static radar. While target detection has been studied in this configuration, parameter estimation has not been investigated to date. To this end, we first develop the signal model for this configuration and derive the CRB for target parameters. The dependence of CRB on system parameters such as SNR, number of snapshots, number of IRS elements and their weights is brought forward through extensive simulations. This study can enable a designer to customize the system parameters to meet the requirements. It also serves as a benchmark for parameter estimation techniques developed for this configuration.
I Introduction
Radar has historically been a vital sensing technology capable of active and passive detection, target tracking, high-resolution imaging, and, increasingly, sensing in Non-Line-of-Sight (NLoS) environments [16, 23, 13, 19]. Conventional mono-static radars usually perform well when targets are within their Line-of-Sight (LoS), enabling precise measurement of range, Doppler shifts, and directions. However, when obstacles block the LoS or targets exhibit stealth behaviour, the effectiveness of traditional radar diminishes significantly or results in complete detection failure, driving the innovation of new sensing architectures.
A common method for enabling NLoS radar sensing involves the use of an Intelligent Reflective Surface (IRS). An IRS is made up of passive, adaptable reflecting elements that can steer incident electromagnetic waves, creating controllable pathways beyond the LoS when positioned appropriately. This IRS-assisted radar technology is well suited in defence, automotive sensing, and modern wireless systems to expand the field-of-view (FOV) and improve NLoS target detection capabilities [22, 20, 4]. Meanwhile, the presence of an IRS introduces multiple possible signal paths from the target to the radar. Specifically, deploying an IRS results in four possible propagation paths [17]:
(P1) $\text{Radar}\rightarrow\text{Target}\rightarrow\text{Radar},$
(P2) $\text{Radar}\rightarrow\text{IRS}\rightarrow\text{Target}\rightarrow\text{IRS}\rightarrow\text{Radar},$
(P3) $\text{Radar}\rightarrow\text{Target}\rightarrow\text{IRS}\rightarrow\text{Radar},$
(P4) $\text{Radar}\rightarrow\text{IRS}\rightarrow\text{Target}\rightarrow\text{Radar}.$
Most IRS-assisted radar studies typically focus on the dominant LoS component (P1) and the IRS-enabled NLoS component (P2) [5]. NLoS sensing via (P2) has been widely studied, including the derivation of Cramér–Rao Lower Bounds (CRLB) for selected target parameters [14]. Despite the beamforming gain yielded by IRS weights in this configuration, it cannot fundamentally mitigate the severe path loss inherent to the four-hop channel, due to the compounded attenuation over the radar–IRS and IRS–target links [12]. This configuration is therefore not suitable for sensing low radar cross-section (RCS) targets and stealth targets.
To mitigate performance degradation due to path loss and scattering from stealth targets, bi-static and multi-static radar configurations are considered. Unlike mono-static radar, in which the transmitter (Tx) and receiver (Rx) are co-located, bi-static and multi-static radars spatially separate the Tx and Rx (s), thereby forming a distributed sensing architecture [10, 18, 6]. This separation enables the receiver to observe the target from a viewpoint that is different from the LoS path between the transmitter and target. In multi-static reception, multiple spatially separated receivers observe the target returns, which can improve detection probability and tracking performance due to spatial diversity [21, 1]. Such configurations are particularly useful for detecting low-observable (stealth) targets and for sensing space objects, making them attractive for modern defence and space applications [15, 7]. Despite these advantages, they introduce important practical limitations, most notably stringent time and phase synchronization requirements between the Tx and distributed Rx (s), along with increased hardware cost and maintenance overhead compared to the mono-static radar.
Upon investigating these trade-offs, we observe the need for a configuration that (i) reduces the effective path loss relative to the double-IRS (P2) route, and (ii) avoids inter-node time synchronization errors inherent in bi-static/multi-static radars. In this work, we focus on the third path (P3), i.e., Radar $\rightarrow$ Target $\rightarrow$ IRS $\rightarrow$ Radar, which is often overlooked but is extremely useful. As will be shown, this path is analogous to a bi-static/multi-static observation geometry because the IRS provides a spatially displaced “listening” viewpoint relative to the radar, yet it remains a passive structure. This motivates a bi-static-like configuration based on IRS-assisted radar. This configuration has been recently presented in [3] for stealth target detection. However, target-parameter estimation within this IRS-assisted framework (P3) has not been studied in the literature.
In this work, the IRS-assisted radar configuration is inspired by the bi-static radar configuration. The position of the IRS panel will coincide with the position of the bi-static receiver which is spatially away from the mono-static radar transceiver. This position is determined by the desired FOV, which is the intersection of the transmitter and the IRS beam patterns. In addition to this requirement, this configuration requires a line-of-sight path between the IRS and the radar transceiver for the IRS to redirect the target’s response towards the radar. A dedicated receive chain can process the signal received from this channel. The sensing chain therefore leverages (P3), mimicking certain advantages of bi-static/multi-static radar while maintaining a mono-static radar front-end, thereby eliminating explicit time synchronization errors between Tx and Rx. Moreover, employing an IRS panel to realize a bi-static-like sensing geometry in challenging scenarios (e.g., stealth target detection) introduces a novel paradigm: from a geometric sensing standpoint, the IRS-assisted propagation path can be interpreted as effectively translating the sensing aperture to the IRS location and subsequently mapping the estimated target coordinates back into the original radar reference frame.
The contributions of the paper are summarized as follows:
  • The IRS-assisted radar configuration is presented in comparison to bi-static/multi-static radar and existing IRS-assisted radars. The associated signal model is systematically developed and reformulated into a parametrized model suitable for the Cramér-Rao Bound(CRB) derivation.
  • We derive CRB for the target angle parameters (azimuth and elevation) under the proposed signal model.
  • Extensive simulation studies are performed to validate the performance of the IRS-assisted radar configuration under different geometric and channel conditions.
The paper is organized as follows: Section II explains in detail the identified IRS-assisted radar configuration. Section III provides the detailed mathematical derivation of the signal model and the target parameters. Based on these target parameters, the theoretical CRB limits are derived in Section IV. Section V discusses the numerical results in relation to the theoretical bounds, and Section VI concludes the paper.
II IRS-Assisted Radar Setup
Bi-static radar separates the transmitter and receiver, forming a signal path with the target as shown in Fig. 1(a)(i), unlike mono-static radar’s setup . This concept generalizes to multi-static radar in Fig. 1(a)(iii), which uses multiple spatially dispersed transmitters and receivers to gain spatial diversity. Multi-static systems observe targets from several angles, enhancing the detection of complex or stealth targets whose Radar Cross Section (RCS) varies with aspect angle. However, effectively exploiting these advantages requires careful spatial arrangement of the receivers, as discussed in [9].
The analogous IRS-assisted radar setup replaces traditional bi-/multi-static receiver nodes with low-cost IRS panels, as shown in Fig. 1(a)(ii) and Fig. 1(a)(iv) that redirects the scattered signals toward nearby radar transceivers [3]. By carefully designing the IRS element weights, the system achieves sufficient beamforming gain to offset longer propagation paths. As shown in Section V, the parameter estimation performance depends on the number of IRS elements with appropriate weights. Fixed IRS placement simplifies radar-IRS channel estimation, and facilitates circumventing synchronization issues typically observed in conventional bi-/multi-static radars.
Amongst the four IRS-assisted configurations discussed in [2] ((P1)-(P4)), the considered IRS-assisted configuration caters to (P1) and (P3) signal paths. This enables the detection of both conventional and stealth targets. Unlike prior IRS-assisted radar work [8], which places the IRS near the radar or target to boost signal strength and relies on a 4-path model (P2) [14], the proposed design positions the IRS panel away from the radar and operates it in simplex mode so that the target response follows the Target → IRS → Radar path. This yields an effective three-path model (P3), reduces path loss compared to conventional four-path IRS-assisted systems, and functionally emulates a bi-static radar configuration. We therefore address this configuration of interest as “IRS-static radar” in the rest of the paper.
((a))
((b))
((c))
Figure 1: Comparison of radar configurations (left), IRS-static radar setup (center), and IRS-static radar field-of-view (FOV) (right).
III Signal Model
Consider a scenario with $K$ targets, with the radar located at $\mathrm{\textbf{r}}_{\mathrm{R}}=[0,0,0]^{T}$ of the coordinate system. Let the $k$th target parameters-namely range, azimuth angle, and elevation angle - be given by $\textbf{r}_{k}$, $\theta^{(\mathrm{az})}_{k}$, and $\theta^{(\mathrm{el})}_{k}$, respectively. Without loss of generality, we consider a 2-D IRS panel, with $L=L_{h}L_{v}$ elements, where $L_{h}$ and $L_{v}$ denote the number of elements along the horizontal and vertical dimensions, respectively, located at $\mathbf{r}_{\mathrm{I}}$ in the radar coordinate system. Let the $k$th target subtend azimuth and elevation angles $\psi^{(\mathrm{az})}_{k}$ and $\psi^{(\mathrm{el})}_{k}$, respectively, as seen from the IRS coordinate system, illustrated in Fig. 1(b). We assume a phased-array radar with $M_{T}=M_{R}$ transceivers co-located at $\mathbf{r}_{\mathrm{R}}$.
When the beamformed transmit signal $\mathbf{s}(t)\in\mathbb{C}^{M_{T}\times 1}$ interacts with the $k$th target, we focus on two scattered components of interest: (i) the monostatic return, and (ii) the component that propagates toward the IRS panel. Denoting the channel between the radar and the $k$th target by $\mathbf{g}_{k}\in\mathbb{C}^{1\times M_{T}}$, the monostatic (ms) component can be expressed as
$\mathbf{y}^{\mathrm{ms}}(t,t_{l})=\sum_{k=1}^{K}\alpha^{\mathrm{ms}}_{k}\mathbf{g}_{k}\mathbf{g}_{k}^{H}\mathbf{s}(t-\tau^{\prime}_{k,l}),$ (1)
where H denotes the conjugate transpose and the channel $\mathbf{g}_{k}$ is assumed reciprocal. Here, $\alpha^{\mathrm{ms}}_{k}$ denotes the monostatic target response strength, $t$ and $t_{l}$ represent the fast-time and slow-time variables, respectively, and $\tau^{\prime}_{k,l}=2(\|\mathbf{r}_{k}-\mathbf{r}_{\mathrm{R}}\|_{2})/c$ is the two-way path delay between the radar and the $k$th target located at $\mathbf{r}_{k}$.
Next, we define the channel between the $k$th target and the IRS panel as $\mathbf{h}_{k}\in\mathbb{C}^{L_{h}L_{v}\times 1}$, and the channel between the IRS panel and the phased-array transceiver as $\mathbf{F}\in\mathbb{C}^{M_{R}\times L}$. The signal received at the radar via the IRS path (denoted as “is”) is given by
$\mathbf{y}^{\mathrm{is}}(t,t_{l})=\sum_{k=1}^{K}\alpha^{\mathrm{is}}_{k}\mathbf{F}\mathbf{D}\mathbf{h}_{k}{\mathbf{g}_{k}}^{H}\mathbf{s}(t-\tau_{k,l}).$ (2)
Let $\alpha^{\mathrm{is}}_{k}$ denote the response strength associated with the IRS-reflected path and the diagonal matrix $\mathbf{D}=\mathrm{diag}(d_{1},\dots,d_{l},\dots,d_{L_{h}L_{v}})\in\mathbb{C}^{L_{h}L_{v}\times L_{h}L_{v}}$ contains the reflection coefficients of all IRS elements. Each coefficient is expressed as $d_{l}=\beta_{l}e^{-j\rho_{l}},$ where $|\beta_{l}|=1$ denotes the unit-modulus amplitude, and $\rho_{l}$ is the phase shift induced by the $l$-th reflecting element of the IRS, for all $1\leq l\leq L_{h}L_{v}$. It is important to note that the total path delay $\tau_{k,l}>\tau^{\prime}_{k,l}$, and is given by
$\tau_{k,l}=\frac{\|\mathbf{r}_{k}-\mathbf{r_{\mathrm{R}}}\|_{2}+\|\mathbf{r}_{k}-\mathbf{r_{\mathrm{I}}}\|_{2}+\|\mathbf{r_{\mathrm{R}}}-\mathbf{r_{\mathrm{I}}}\|_{2}}{c}.$ (3)
Accordingly, the overall received signal at the transceiver can be written as
$\mathbf{y}(t,t_{l})=\mathbf{y}^{\mathrm{ms}}(t,t_{l})+\mathbf{y}^{\mathrm{is}}(t,t_{l})+\mathbf{v}(t,t_{l}),$ (4)
where $\mathbf{v}(t,t_{l})$ is an additive white Gaussian noise (AWGN) vector. For conventional targets, $\alpha^{\mathrm{ms}}_{k}$ is typically large; hence, returns from both paths must, in principle, be processed. However, because the path delays differ between the monostatic path and the path via the IRS, the signals must be processed independently. This can be accomplished by beam-space processing or by having a dedicated receive array looking towards the IRS panel. Since parameter extraction from conventional targets using the monostatic path is well studied, we focus our study on the parameter estimation from the IRS path. The signal received by this array from only the IRS path is given by:
$\small\begin{split}\mathbf{\overline{y}}(nT_{s},t_{l})&=\sum_{k=1}^{K}\alpha^{\mathrm{is}}_{k}\mathbf{F}\mathbf{D}\mathbf{h}_{k}\mathbf{g}_{k}^{H}\mathbf{s}(nT_{s}-\tau_{k,l})+\mathbf{v}(nT_{s},t_{l}).\end{split}$ (5)
It is observed that the IRS weights in $\mathbf{D}$ influence the received signal in (5). We next investigate the design of the IRS coefficients and the reconstruction of the underlying signal model in a manner that facilitates estimation of the target parameters in the following subsections.
III-A Design of IRS Weights
To achieve performance analogous to a bi-static/multi-static radar, the IRS-static radar must incorporate the effective field-of-view (FOV) of the corresponding distributed configuration into its design. From the IRS-static radar depicted in Fig. 1(c), we note that the bistatic receiver is replaced by the IRS panel. To realize the desired FOV, the bi-static radar receiver would be oriented toward the direction $\big(\psi^{\left(\mathrm{az}\right)}_{0},\psi^{\left(\mathrm{el}\right)}_{0}\big)$. For the IRS-static radar, the IRS weights must be configured such that signals received from this angular region are redirected toward the radar receiver, which is located at an angle $\big(\psi^{\left(\mathrm{az}\right)}_{\mathrm{R}},\psi^{\left(\mathrm{el}\right)}_{\mathrm{R}}\big)$ relative to the IRS panel. The coherent signal arrival at the receive array serves as the design objective for IRS weights. Defining the diagonal matrices $\mathbf{A}_{0}=\mathrm{diag}\{\mathbf{a}(\psi^{\left(\mathrm{az}\right)}_{0},\psi^{\left(\mathrm{el}\right)}_{0})\}$ and $\mathbf{A}_{\mathrm{R}}=\mathrm{diag}\{\mathbf{a}(\psi^{\left(\mathrm{az}\right)}_{\mathrm{R}},\psi^{\left(\mathrm{el}\right)}_{\mathrm{R}})\}$, the diagonal IRS weight matrix $\mathbf{D}$ should satisfy
$\mathbf{A}_{0}\mathbf{D}\mathbf{A}_{\mathrm{R}}=\mathbf{I},$ (6)
where $\mathbf{I}$ is the identity matrix. The resulting solution is
$\mathbf{D}=\mathbf{A}_{0}^{*}\mathbf{A}_{\mathrm{R}}^{*}.$ (7)
For targets in the vicinity of $\big(\psi^{\left(\mathrm{az}\right)}_{0},\psi^{\left(\mathrm{el}\right)}_{0}\big)$, the IRS panel redirects the response back to the radar with some beamforming gain, attaining maximum gain when the target is located exactly in the direction $\big(\psi^{\left(\mathrm{az}\right)}_{0},\psi^{\left(\mathrm{el}\right)}_{0}\big)$. Although this design does not incorporate the desired beamwidth that determines the FOV, we employ this for the target-parameter estimation CRB analysis. The performance of this IRS weight design is analyzed in Section V.
III-B Reconstruction of Signal Model
Stacking the received signal in (5) over $C$ chirps, we obtain the radar data cube $\mathbf{\overline{Y}}\in\mathbb{C}^{M\times N\times C}$, where $N$ denotes the number of fast-time samples. Performing the discrete Fourier transform (DFT) along the slow-time dimension, we to obtain $M_{R}$ Range-Doppler (RD) maps, $\mathbf{\overline{Y}}_{\mathrm{RD}}=\mathcal{F}_{C}\{\mathbf{\overline{Y}}\}$, where $\mathcal{F}_{C}\{.\}$ is a $C$-point DFT operator. Due to the superposition of the $K$ target responses, $\mathbf{\overline{Y}}_{\mathrm{RD}}=\sum\limits_{k=1}^{K}\mathbf{\overline{Y}}^{k}_{\mathrm{RD}}$, where $\mathbf{\overline{Y}}^{k}_{\mathrm{RD}}$ denotes the RD contribution due to the $k$th target alone. Using either constant false-alarm-rate (CFAR) detector (or its variants) [9] or technique developed in [3], one can obtain $K_{P}\leq K$ RD bins corresponding to true detections; with some bins potentially containing multiple targets. Therefore, the total number of targets $K=\sum_{p=1}^{K_{P}}P_{k}$, where $P_{k}$ denotes the number of targets present in the $k$th detected RD bin, $(\hat{R}_{k},\hat{v}_{k})$. Note that $\hat{R}_{k}=\frac{c\tau_{k,l}|_{l=0}}{2}$ corresponds to the round-trip distance via the IRS, while $\hat{v}_{k}$ denotes the relative bi-static velocity of the target with respect to the radar and the IRS.
The observation vector for the $k$th RD bin across the $M_{R}$ receivers is given by
$\mathbf{\overline{y}}_{k}=\mathbf{FD}\sum_{p=1}^{P_{k}}\mathbf{h}_{p}\mathbf{g}_{p}^{H}\mathbf{s}+\mathbf{\bar{v}}_{k}.$ (8)
We denote the LoS channel between the radar and the $p$th target as $\mathbf{g}_{p}=\alpha_{p}^{(g)}\mathbf{a}_{M_{T}}(\theta^{(\mathrm{az})}_{p},\theta^{(\mathrm{el})}_{p})$, where $\mathbf{a}_{M_{T}}(\theta^{(\mathrm{az})}_{p},\theta^{(\mathrm{el})}_{p})$ is the steering vector toward the target as seen from the transmit array. Similarly, we write $\mathbf{h}_{p}=\alpha_{p}^{(h)}\mathbf{a}_{L_{h}L_{v}}\left(\psi^{(\mathrm{az})}_{p},\psi^{(\mathrm{el})}_{p}\right)$, where $\mathbf{a}_{L_{h}L_{v}}\left(\psi^{(\mathrm{az})},\psi^{(\mathrm{el})}\right)=\mathbf{a}_{L_{v}}\left(\psi^{(\mathrm{az})}\right)\otimes\mathbf{a}_{L_{h}}\left(\psi^{(\mathrm{el})}\right)$ is the steering vector toward the target as observed from the IRS panel, and $\otimes$ denotes the Kronecker product. Defining $\alpha_{p}=\alpha_{p}^{(h)}\alpha_{p}^{(g)}$, (8) can be rewritten as
$\mathbf{\overline{y}}_{k}=\mathbf{FD}\sum_{p=1}^{P_{k}}\alpha_{p}\mathbf{a}_{L_{h}L_{v}}\left(\psi_{p}^{\left(\mathrm{az}\right)},\psi^{\left(\mathrm{el}\right)}_{p}\right)\\ \times\mathbf{a}_{M_{T}}\left(\theta^{\left(\mathrm{az}\right)}_{p},\theta^{\left(\mathrm{el}\right)}_{p}\right)^{H}\mathbf{\overline{s}}+\mathbf{\bar{v}}_{k}.$ (9)
With reference to Fig. 1(b) for the detection obtained at ($\hat{R}_{k}$$\hat{v}_{k}$), the parameters $\theta^{\left(\mathrm{az}\right)}_{k},\theta^{\left(\mathrm{el}\right)}_{k},\psi^{\left(\mathrm{az}\right)}_{k},\psi^{\left(\mathrm{el}\right)}_{k},\forall\ k\in[1,K]$ are to be estimated. Importantly, the 3-D target position can be determined using only a subset of these parameters. Let $R_{k,\mathrm{R}}$ and $R_{k,\mathrm{I}}$ denote the distances from the $k$th target to the radar and to the IRS, respectively. Then the estimated range inferred from the path delay satisfies
$\hat{R}_{k}=R_{k,\mathrm{R}}+R_{k,\mathrm{I}}+R_{\mathrm{I},\mathrm{R}},$ (10)
where the distance between IRS and the radar $R_{\mathrm{I},\mathrm{R}}$ is assumed known. It follows that estimating $\psi^{(\mathrm{az})}_{k}$ and $\psi^{(\mathrm{el})}_{k}$, together with $R_{k,\mathrm{I}}$, yields the $k$th target position $\mathbf{r}_{k}^{\prime}=[x_{k}^{\prime},y_{k}^{\prime},z_{k}^{\prime}]^{T}$ with respect to the IRS coordinate system. The target position in the radar coordinate system can then be obtained via
$\mathbf{r}_{k}=\mathbf{r}_{k}^{\prime}+\mathbf{r}_{\mathrm{I}}.$ (11)
Alternatively, the target position can be obtained directly by estimating $\theta^{\left(az\right)}_{k}$ and $\theta^{\left(el\right)}_{k}$ along with $R_{k,\mathrm{R}}$ from $\hat{R}_{k}$.
It is evident from (9) that the received signal corresponds to a point in 3-D space that is defined by the intersection of $\mathbf{a}_{L_{v}L_{h}}\left(\psi_{p}^{\left(\mathrm{az}\right)},\psi_{p}^{\left(\mathrm{el}\right)}\right)$ and $\mathbf{a}_{M_{T}}\left(\theta_{p}^{\left(\mathrm{az}\right)}\right)$, further constrained by the round-trip delay. (9) can be compactly represented as:
$\overline{\textbf{y}}_{k}=\textbf{F}\textbf{D}\textbf{A}\overline{\textbf{x}}+\overline{\textbf{v}}_{k},$ (12)
where $\overline{\textbf{x}}=\sum^{P_{k}}_{p=1}\alpha_{p}\textbf{a}_{MT}(\theta^{az}_{p},\theta^{el}_{p})\textbf{s}$ and the dictionary matrix $\mathbf{A}\in\mathbb{C}^{L\times G_{R}}$ corresponds to the angle-of-arrival space around the IRS such that
$\mathbf{A}=\big[\mathbf{a}(\psi_{1}),\dots,\mathbf{a}(\psi_{G_{R}})\big],$ (13)
where the angular grid points $\psi_{1},...,\psi_{G_{R}}]$ are expressed as ${\Psi}=\{\psi_{z}:\psi_{z}=\left(\psi_{z}^{(\mathrm{az})},\psi_{z}^{(\mathrm{el})}\right),\psi_{z}^{(\mathrm{az})}\in\left[-\pi/2,\pi/2\right],\psi_{z}^{(\mathrm{el})}\in\left[0,\pi/2\right]\ 1\leq z\leq G_{R}\}$. Based on the target field of view with respect to the IRS, the angular region (and hence the number of grid points, $G_{R}$) can be reduced. In the following section, Cramer-Rao Bounds are derived for the signal model presented above.
IV Cramér-Rao Bounds
The target parameters to be estimated from all detected observation vectors $\overline{\mathbf{y}}_{k}$ are
$\boldsymbol{\zeta}$ $=[\zeta_{1},\zeta_{2},\ldots,\zeta_{K}],$ (14)
where, $\boldsymbol{\zeta}_{i}=[\psi_{i}^{(\mathrm{az})},\,\psi_{i}^{(\mathrm{el})}]^{T}$is the target parameters correponding to $k$-th target and $\mathbf{\zeta}\in\mathbb{R}^{2\times K}$ is the target parameters for all $K$ targets as seen from the IRS. Our objective is to characterize the fundamental error bound, i.e., the minimum achievable error variance, for estimating these angle parameters under the proposed signal model using the Cramér–Rao Bound (CRB) framework.
Several approaches exist for deriving CRB depending on the assumptions and the structure of the signal model [11]. In other IRS-assisted radar configurations, CRB has been derived for signal waveform estimation in [5] and for target parameters estimation in [17]. In this work, we derive CRB for target parameter estimation assuming deterministic transmit signal waveform in (P3) IRS-assisted configuration. We develop CRB for IRS-based $\zeta_{i},i\in[1,K]$, which is equivalent to the study w.r.t. the radar because of the positional equivalence observed in (11).
For CRLB analysis, the proposed signal model in (12) is rearranged as
$\overline{\textbf{y}}_{k}=\overline{\textbf{A}}\overline{\textbf{x}}+\overline{\textbf{v}},$ (15)
where $\overline{\textbf{A}}=\textbf{F}\textbf{D}\textbf{A}$. This compact form of the signal model ensures that only the target parameters as seen from the IRS, i.e., $(\psi^{\mathrm{(az)}},\psi^{\mathrm{(el)}})$, are explicit while the rest are embedded within $\overline{\mathbf{A}}$.
The Fisher Information Matrix (FIM) associated with the above signal model and the parameter vector of interest is given by
$\small\mathbf{J}=\begin{bmatrix}\mathbf{J}_{\psi^{(\mathrm{az})},\psi^{(\mathrm{az})}}&\mathbf{J}_{\psi^{(\mathrm{az})},\psi^{(\mathrm{el})}}\\ \mathbf{J}_{\psi^{(\mathrm{el})},\psi^{(\mathrm{az})}}&\mathbf{J}_{\psi^{(\mathrm{el})},\psi^{(\mathrm{el})}}\end{bmatrix}$. (16)
To obtain $\mathbf{J}$, we first define the vector derivative for each column of $\overline{\textbf{A}}$ with respect to target parameters as
$\mathbf{Z}_{i}=\left[\frac{\partial\mathbf{\overline{a}}(\boldsymbol{\zeta}_{i})}{\partial\psi_{i}^{(\mathrm{az})}}\quad\frac{\partial\mathbf{\overline{a}}(\boldsymbol{\zeta}_{i})}{\partial\psi_{i}^{(\mathrm{el})}}\right]\quad\in\mathbb{C}^{M_{R}\times 2},i=1,\cdots,K,$ (17)
Stacking these derivatives of all target parameters for $K$ targets, we get
$\mathbf{Z}=\left[\mathbf{Z}_{1}\quad\mathbf{Z}_{2}\quad\cdots\quad\mathbf{Z}_{K}\right]\quad\in\mathbb{C}^{M_{R}\times 2K},$ (18)
Defining $\textbf{H}=\textbf{Z}^{H}\textbf{P}_{\frac{1}{\overline{\textbf{A}}}}\textbf{Z},$, where $\textbf{P}_{\frac{1}{\overline{\textbf{A}}}}=\textbf{I}-\overline{\textbf{A}}(\overline{\textbf{A}}^{H}\overline{\textbf{A}})^{-1}\overline{\textbf{A}}^{H}$ is the orthogonal projection matrix of $\overline{\textbf{A}}$, (16) can be compactly expressed as [11]:
$\textbf{J = }\frac{2N}{\sigma_{w}^{2}}\,\mathrm{Re}\!\left[\mathbf{H}\odot\left(\mathbf{S}_{\mathbf{f}}^{T}\otimes\mathbf{1}_{2\times 2}\right)\right],$ (19)
where $\sigma_{w}^{2}$, $N$, $S_{f}$ and $\mathbf{1}_{2\times 2}$ denote the noise variance, the number of snapshots, the source signal powers and $2\times 2$ matrix of ones, respectively. $\odot$ is the Hadamard product operator. The resulting CRB expression is given by
$\textbf{C}_{CR}=\textbf{J}^{-1}=\begin{bmatrix}\mathbf{C}_{\psi^{\mathrm{(az)}},\psi^{\mathrm{(az)}}}&\mathbf{C}_{\psi^{\mathrm{(az)}},\psi^{\mathrm{(el)}}}\\ \mathbf{C}_{\psi^{\mathrm{(el)}},\psi^{\mathrm{(az)}}}&\mathbf{C}_{\psi^{\mathrm{(el)}},\psi^{\mathrm{(el)}}}\end{bmatrix}$. (20)
The $\mathbf{C}_{CR}$ contains the variance corresponding to the azimuth and elevation of each target. Consequently, the Root-Mean-Square Error (RMSE) for azimuth and elevation estimates are obtained as
$RMSE(\psi^{(\mathrm{az})})=\sqrt{\textbf{C}_{\psi^{(\mathrm{az})},\psi^{(\mathrm{az})}}},$ (21)
$RMSE(\psi^{(\mathrm{el})})=\sqrt{\textbf{C}^{\psi^{(\mathrm{el})},\psi^{(\mathrm{el})}}}.$ (22)
Although the derived CRB resembles the standard form of a linear signal model, the critical contribution of the IRS path to the performance of the IRS-static radar configuration is embedded within the CRB. Besides the dependence of CRB on the snapshots ($N$) and the SNR $(\sigma^{2}_{w},S_{f})$, the matrix $H$ embeds the system setup in relation with target positions as seen from the IRS. Since the considered IRS-static configuration is new, we employ the derived CRB to analyse the sensitivity/robustness of this configuration across various target and system scenarios for the target parameter estimation in the subsequent section.
V Results
We evaluate the performance of the proposed IRS-static radar system using a simulated 24  GHz sensing configuration. The radar transmitter (Tx) and receiver (Rx) arrays are co-located at the origin of the coordinate system. The transmitter and receiver array apertures are uniform linear arrays (ULA) composing $M_{T}=M_{R}=50$ elements placed along the $x$-axis, with half-wavelength spacing $d_{\mathrm{radar}}=\lambda/2=6.25$ mm. Hence, the $m$-th receive sensor is located at $\bigl((m-1)d_{\mathrm{radar}},\,0,\,0\bigr)$.
A single IRS panel is positioned at $\textbf{r}_{I}=\bigl[-50,\,100,\,0\bigr]^{T}$ m and realized as a $20\times 18$ rectangular grid of reflecting elements, resulting in $L=360$ controllable units. The IRS elements are arranged in the yyzz plane with half-wavelength spacing along both axes. The FOV of the IRS-static radar is the intersection of the transmit beam and the IRS look direction, as shown in Fig. 1(c). Accordingly, the IRS weights are computed using (7) so that the FOV is steered to the point $(\Psi^{(\mathrm{az})}_{0},\Psi^{(\mathrm{el})}_{0})$, and the reflected signal is directed toward the radar receive array at $(\Psi^{(\mathrm{az})}_{\mathrm{R}},\Psi^{(\mathrm{el})}_{\mathrm{R}})$. Being agnostic to any parameter estimation technique, we study the CRB against various system and environment parameters.
((a))
((b))
Figure 2: RMSE performance versus angular mismatch between the target and IRS lookup direction in (a) Azimuth and (b) Elevation.
[…]
It is interesting to note that the RMSE curve for azimuth angle deviation exhibits a narrow beam around $\Psi^{\mathrm{(az)}}_{0}$ as seen in Fig. 2(a), compared to that observed for elevation angle around $\Psi^{\mathrm{(el)}}_{0}$ in Fig. 2(b). This is because of the large effective aperture due to the IRS elements and the radar receive array in the x-y plane. For the elevation angle estimation, the wide RMSE curve around $\Psi^{\mathrm{(el)}}_{0}$ in Fig. 2(b) can be attributed to the reduced effective aperture of IRS elements present along the z-plane. Furthermore, the inverted sidelobes do not exhibit significantly high RMSE compared to the mainlobe, indicating target response from outside the FOV will receive a reasonable gain unless the target lies exactly on the beams’ nulls. From this observation, it can be inferred that the RMSE performance, the FOV, and the sidelobe performance of the IRS-static radar can be altered through appropriate design of the IRS weights. Conversely, the required FOV along the azimuth and elevation angles have to be critical design parameters for IRS weights in this context.
((a))
((b))
Figure 3: RMSE performance versus (a) SNR and (b) Number of snapshots.
[…]
A monotonous CRB decrease is observed in Fig. 3(a) with increasing SNR. For $\Psi_{0}=\Psi_{k}$, the RMSE error for azimuth angle is distinctly lesser than the elevation angle over the entire SNR range due to the larger effective aperture along the x-y dimension over the z-dimension as discussed earlier. Similarly, Fig. 3(b) examines the dependence of azimuth and elevation RMSE on the number of snapshots. Both error metrics decrease as the number of snapshots increases from the order of hundreds to the order of thousands, with progressively diminishing gains thereafter, once the covariance estimate becomes more stable.
When we specifically observe the plots corresponding to deviation of $0.5^{\circ}$, i.e., $\Psi^{(\mathrm{az})}_{0}-\Psi^{(\mathrm{az})}_{k}=0.5^{\circ}$ and $\Psi^{(\mathrm{el})}_{0}-\Psi^{(\mathrm{el})}_{k}=0.5^{\circ}$ (azimuth–pink and elevation–black), one can note that the RMSE is higher than the condition when $\Psi^{(\mathrm{az})}_{0}=\Psi^{(\mathrm{az})}_{k}$ and $\Psi^{(\mathrm{el})}_{0}=\Psi^{(\mathrm{el})}_{k}$ because the target is off the IRS look direction. The azimuth RMSE is lower than elevation RMSE, similar to the previous case, as both these deviations lie in the main lobe corresponding to Fig. 2(a) and Fig. 2(b). For the case where $\Psi^{(\mathrm{az})}_{0}-\Psi^{(\mathrm{az})}_{k}=2.5^{\circ}$ and $\Psi^{(\mathrm{el})}_{0}-\Psi^{(\mathrm{el})}_{k}=2.5^{\circ}$, the corresponding plots (azimuth–green and elevation–gold) in Fig. 3(a) and Fig. 3(b) reveal that the RMSE curves are higher than those of the $0^{\circ}$ and $0.5^{\circ}$ cases. This is due to the deviation from the look-up direction being larger compared to the $0.5^{\circ}$ case, resulting in larger loss in directional gain at the IRS. Furthermore, it is interesting to note that the elevation RMSE is less than the azimuth RMSE. This is because the target azimuth angle lies in the first sidelobe of the pattern observed in Fig. 2(a) while the target elevation angle still resides in the mainlobe pattern shown in Fig. 2(b).
Figure 4: RMSE performance as a function of the number of IRS elements.
Figure 4 illustrates the dependence of azimuth and elevation RMSE on the number of IRS elements, $L=L_{h}L_{v}$. In this study, as the number of IRS elements increases, we maintain a square IRS panel with $L_{h}=L_{v}\in[2,40]$. As $L$ increases from $4$ to $1600$, the azimuth angle and elevation angle RMSE decrease, proving the role of aperture gain in IRS-assisted sensing. Larger IRS apertures enhance the effective array gain and enable more accurate phase synthesis, thereby strengthening the reflected signals and improving angular resolution. The RMSE reduction is most pronounced in the range $4\leq L\leq 400$, after which the curves tend to saturate, indicating diminishing returns with further aperture enlargement. The RMSE of azimuth is lower than that of elevation, following the similar trend observed in Figs. 3(a) and  3(b). Overall, increasing the number of IRS elements is particularly beneficial in the low- to mid-aperture regime, whereas for larger apertures the design must balance incremental performance gains against the FOV and sidelobe attenuation.
Furthermore, for the design of IRS weights as illustrated in (7), it is to be noted that the beamwidth decreases with increasing aperture length, thereby affecting the FOV. We remark that the design of IRS weights using well-constructed optimization problem takes the center stage in the performance of the IRS-static radar configuration for the desired FOV.
VI Conclusion
In this work, a three-hop IRS-assisted radar configuration, termed as IRS-static radar, is considered for study. It is shown that this configuration mimics a bi-static radar with time synchronization alleviated with IRS panel redirecting the scattered signal back to the mono-static radar. A signal model is formulated and compactly represented for this configuration to derive the CRB for target parameters. The dependence of CRB on the system parameters such as the SNR, number of snapshots, number of IRS elements and their weights is brought forward through extensive simulations. The CRB limit is studied as the target moves away from the IRS look direction, both in azimuth and elevation directions, seeking appropriate design for the FOV. It is therefore inferred that the beamformer design for IRS takes center stage in the performance of the IRS-static radar. This study enables a designer to customize the system parameters as per the system specification and serves as a benchmark for parameter estimation techniques developed for this configuration.